# BOOM: Benchmarking Out-Of-distribution Molecular Property Predictions of Machine Learning Models

### v1

@@ 4.4 Representation
Representation Split HoF Density HOMO LUMO Gap ZPVE $R^{2}$ $\alpha$ $\mu$ $C_{V}$
3D ID 11.09 .0121 .0026 .0030 .0042 .0005 21.7465 .3234 .3690 .1296
OOD 21.76 .0247 .0152 .0137 .0238 .0031 112.7228 .30890 2.2832 .9457
Explicit Bonds ID 15.68 .0092 .0041 .0048 .0058 0.0014 35.68 .8305 .55 .3341
OOD 100.6 .0551 .0192 .0187 .0267 .0129 234.73 .4772 2.3 2.149
SMILES ID 22.86 .0163 .0068 .0088 .0103 .0046 50.297 1.444 .7134 .4923
OOD 99.7253 .1173 .0245 .0267 .0315 .0214 306.14 6.303 2.766 3.0175
@@
Table 3: RMSE of models on OOD and ID tasks as grouped by input representation. The best performing ID and OOD models are highlighted in Black and Blue respectively. The worst performing ID and OOD models are highlighted in Orange and Red respectively.
In our study, 3D models with equivariant and invariant symmetries significantly outperform the SMILES-based models in nearly all tasks. Furthermore, the 3D GNN models like EGNN and IGNN are significantly more parameter-efficient. As we can see in Figure 3, the SMILES-based models, namely the transformer models, perform significantly worse than the 3D and explicit structure models in nearly all tasks. SMILES and explicit bonds are interchangeable representations in that SMILES can be converted into a molecular graph and vice-versa. SMILES-based representations present the same atom and topology information present in graph-like explicit-bonds representation but in a sequence format. This suggests the inductive bias present in the graph-based models improve the model performance over attention-based models, especially for OOD splits. Interestingly, the graph-like models also perform comparatively to the transformer-based models if we discount RT. MolFormer, a SMILES based, model has strong ID performance compared to other models as well.
4.5 Data Augmentation
@@
Figure 8: Performance of the Chemprop MPNN model when various amounts of OOD samples are included in the training. For each property, we separate the 10,000 OOD examples into a hold-out test set (N=5000) and various amounts of the remaining 5000 examples are included during training. The OOD test RMSE is normalized against the validation RMSE of the model trained without any additional OOD examples included during training.
Beyond exploring different model architectures and molecular representations, data generation is a common strategy for improving the generalization capabilities of deep learning models. As a notable example, the authors of the GNoME material dataset showed that the out-of-distribution property prediction performance of a deep learning model can be vastly improved by training on an augmented dataset of millions of generated materials. [36] Similarly, using active learning to iteratively retrain graph neural networks was recently shown to reduce the property prediction error on OOD generated molecules for both density (75% reduction) and solid heat of formation (83% reduction). [8]
Figure 8 investigates improving OOD property prediction by augmenting the QM9 training set ( described in Section 1.3) with extreme-valued molecules from the QM9 OOD test set. We select N =[0, 100, 500, 1000, 2000, 5000] molecules from the original QM9 test set by selecting molecules with properties below and above the 25th and 75th quantiles, respectively.
Across 7 of 8 QM9 tasks, Chemprop’s generalization improves with augmented data (Figure 8). The lack of improvement for QM9 dipole moments (Mu) likely stems from Chemprop’s graph representation lacking 3D electronic structure. Data augmentation consistently yields sizable generalization improvements, even with a small fraction ( 4%) of augmented data. Further improvements may be achievable with more extensive data generation.
4.6 Improving Underperforming Models
Finally, we highlight a significant improvement in OOD performance with ModernBERT among the NLP-style models tested. While all the NLP models tested don’t have any chemistry specific design choices, the improvements proposed in ModernBERT translate to the chemistry domain as well. We highlight the task-specific behavior in Figure 9 for OOD performance. ModernBERT performs similar to other transformer models for the difficult tasks highlighted in Figure 4, but improves significantly for the remaining properties. We see improvements in both intensive and extensive properties. ModernBERT decreases OOD HoF and $C_{v}$ RMSE by more than 58% and 78% respectively over other best performing transformer models. While not in the scope of our current work, understanding the design choices that result in these improvements can inform design choices for future chemistry foundation model design.
Figure 9: ModernBERT outperforms transformer models for OOD tasks. ModernBERT bridges the gap between transformer and GNN-based models on OOD splits, especially for HoF, Density, and $C_{v}$.
5 Discussion
In this work, we have developed BOOM, a standardized benchmark for evaluating the OOD performance of chemical property prediction models. Overall, across all 12 tested model architectures we do not find any model that achieves strong performance on all OOD tasks. As a result, we expect that current property prediction models will struggle to consistently discover molecules with properties that extrapolate beyond known molecules. Nevertheless, given the saturation of the most commonly used chemistry benchmarks, we hope that the results presented here inspire the chemistry community to pursue OOD generalization as the next frontier challenge for further developing molecular property prediction models.
Our findings also highlight promising future directions towards improving OOD generalization. Surprisingly, we found that commonly employed molecular pretraining strategies, such as masked language modeling, often result in a decrease in OOD performance. Our experiments show that developing new pretraining tasks whereby the pretraining task and the downstream property prediction tasks are more closely related results in improved OOD generalization. Fig. 6 consistently shows that OOD performance is only improved by pretraining when the chemical information contained in the pretraining task is related to the downstream property prediction task. Randomly sampling model hyperparameters of the Chemprop GNN was found to significantly improve OOD performance, with very little change in ID performance. This result highlights the need to consider OOD generalization when optimizing the model hyperparameters of chemical prediction models.
While high-inductive bias and 3D models perform well on our current tests, scalability remains a significant issue. Small models can provide strong predictive power, but they do not allow for techniques such as in-context learning and test-time compute that may be available to large scale models. Similarly, 3D models are attractive, but high-quality DFT data is not always available. While 3D molecular data is becoming increasingly more available, it dwarfs in comparison to the billions of molecules used in unsupervised molecular pretraining strategies. In general, as our results with ModernBERT show, transformer-based models can potentially catch up to the small models while enabling greater scalability. Numerical encoding is a concern for LLM-like models and was a significant drawback for RT. Improved post-hoc solutions [37] or modern tokenization techniques [38, 39] will be key in the development of LLM-based predictive models.
Perhaps unsurprisingly, augmenting the training dataset to minimize the distribution shift between the ID and OOD samples was found to consistently improve model generalization. However, the ease of generating a sufficient quantity of high-quality molecular property data to augment the training set is highly task-dependent, which may limit the applicability of this approach.
Notably, we do not find any strategy that universally improves OOD performance across all property prediction tasks. On relatively simple properties (i.e. density and ZPVE), current SOTA property prediction models exhibit strong generalization with very little difference between ID and OOD performance. However, we note that properties that depend on the electronic structure of the molecules (i.e. Dipole moment and HOMO) showed little improvement in OOD performance across all experimental settings. We anticipate that achieving strong OOD generalization on these properties will require larger datasets, in combination with molecular representations that explicitly capture the molecules’ electronic structure.
6 Code Availability
The BOOM benchmark datasets will be made available open-source on Github. This repository contains code for generating consistent OOD property prediction evaluations, as well as the benchmarking results presented in this paper.
7 Acknowledgments
This work was performed under the auspices of the U.S. Department of Energy by Lawrence Livermore National Laboratory under Contract DE-AC52-07NA27344. This manuscript was supported by LDRD 24-SI-008 (LLNL-JRNL-2004980).

### v2

@@ C.6 Tabulated results
Model Type Split HoF Density HOMO LUMO GAP ZPVE $R^{2}>$ $\alpha$ $\mu$ $C_{v}$
ChemBERTa Transformer ID 0.937 0.963 0.838 0.963 0.944 0.992 0.948 0.977 0.649 0.985
OOD 0.339 0.372 0.059 0.021 0.076 0.544 0.063 0.009 0.104 0.126
Chemprop Explicit-Bonds ID 0.963 0.985 0.940 0.987 0.980 0.996 0.967 0.981 0.792 0.990
OOD 0.528 0.901 0.206 0.161 0.078 0.597 0.331 0.332 0.096 0.463
EGNN 3D ID 0.992 0.988 0.929 0.933 0.988 0.981 0.994 0.980 0.979 0.765
OOD 0.869 0.886 0.238 0.074 0.028 0.681 0.400 0.344 0.078 0.424
ET 3D ID 0.977 0.994 0.974 0.994 0.989 1.000 0.975 0.997 0.702 0.999
OOD 0.861 0.953 0.298 0.313 0.103 0.952 0.684 0.673 0.059 0.870
Geoformer 3D ID 0.954 0.991 0.973 0.995 0.989 1.000 0.997 0.997 0.740 0.999
OOD 0.695 0.941 0.243 0.276 0.035 0.831 0.835 0.401 0.034 0.683
GotenNet 3D ID 0.996 0.992 0.926 0.993 0.978 1.000 0.997 0.994 0.929 0.998
OOD 0.946 0.947 0.151 0.335 0.026 0.999 0.998 0.772 0.111 0.990
Graphormer 3D ID 0.985 0.991 0.944 0.989 0.982 1.000 0.971 0.995 0.723 0.997
OOD 0.790 0.926 0.304 0.249 0.272 0.445 0.482 0.351 0.117 0.651
IGNN 3D ID 0.973 0.987 0.922 0.985 0.970 0.994 0.977 0.976 0.812 0.986
OOD 0.852 0.866 0.229 0.544 0.035 0.737 0.424 0.207 0.090 0.353
MACE 3D ID 0.995 0.412 0.373 0.889 0.807 0.996 0.997 0.998 0.867 0.998
OOD 0.846 0.212 0.025 0.184 0.008 0.979 0.858 0.823 0.058 0.990
MLP No Bias ID 0.983 0.091 0.000 0.765 0.708 0.785 0.880 0.976 0.611 0.982
OOD 0.784 0.089 0.017 0.086 0.003 0.142 0.283 0.166 0.093 0.903
MoLFormer Transformer ID 0.984 0.992 0.927 0.985 0.978 0.999 0.970 0.992 0.782 0.995
OOD 0.281 0.604 0.044 0.028 0.097 0.480 0.187 0.153 0.076 0.453
ModernBERT Transformer ID 0.955 0.980 0.880 0.972 0.957 0.999 0.957 0.984 0.685 0.991
OOD 0.691 0.893 0.285 0.230 0.190 0.992 0.731 0.764 0.092 0.918
RT Transformer ID 0.954 0.969 0.568 0.958 0.938 0.981 0.938 0.926 0.472 0.977
OOD 0.180 0.226 0.020 0.017 0.032 0.044 0.096 0.059 0.085 0.084
Random Forest No Bias ID 0.921 0.912 0.874 0.969 0.957 0.999 0.933 0.979 0.679 0.988
OOD 0.147 0.021 0.023 0.013 0.043 0.041 0.084 0.108 0.098 0.011
TGNN Explicit-Bonds ID 0.970 0.961 0.872 0.963 0.869 0.998 0.407 0.986 0.681 0.989
OOD 0.814 0.841 0.268 0.284 0.057 0.990 0.131 0.554 0.074 0.926
@@
Table 5: Mean Batched $R^{2}$ scores of all models on OOD and ID tasks. Best performing ID and OOD models are highlighted in Black and Blue respectively.The worst performing ID and OOD models are highlighted in Orange and Red respectively. The graph-based and hybrid models provide the best scores across nearly all tasks for OOD and ID splits.
Model Type Split HoF Density HOMO LUMO GAP ZPVE $R^{2}>$ $\alpha$ $\mu$ $C_{v}$
ChemBERTa Transformer ID 0.002 0.005 0.002 0.001 0.002 0.003 0.001 0.001 0.002 0.002
OOD 0.040 0.050 0.013 0.004 0.014 0.062 0.023 0.007 0.003 0.103
Chemprop Explicit-Bonds ID 0.001 0.001 0.001 0.001 0.001 0.001 0.001 0.002 0.002 0.001
OOD 0.011 0.005 0.008 0.033 0.016 0.059 0.007 0.016 0.000 0.027
EGNN 3D ID 0.003 0.003 0.005 0.009 0.003 0.006 0.002 0.012 0.004 0.025
OOD 0.037 0.005 0.014 0.009 0.005 0.033 0.007 0.128 0.001 0.054
ET 3D ID 0.007 0.002 0.001 0.001 0.000 0.000 0.025 0.002 0.354 0.000
OOD 0.037 0.012 0.022 0.011 0.011 0.010 0.173 0.039 0.011 0.023
Geoformer 3D ID 0.005 0.001 0.001 0.000 0.000 0.000 0.001 0.000 0.010 0.000
OOD 0.101 0.003 0.021 0.025 0.040 0.012 0.074 0.059 0.003 0.018
GotenNet 3D ID 0.001 0.000 0.003 0.001 0.001 0.000 0.000 0.003 0.005 0.000
OOD 0.011 0.005 0.019 0.031 0.014 0.000 0.000 0.011 0.008 0.001
Graphormer 3D ID 0.004 0.003 0.002 0.001 0.001 0.001 0.001 0.001 0.003 0.000
OOD 0.053 0.009 0.107 0.180 0.164 0.155 0.062 0.020 0.024 0.053
IGNN 3D ID 0.006 0.001 0.004 0.001 0.002 0.007 0.003 0.012 0.016 0.006
OOD 0.005 0.019 0.015 0.354 0.001 0.024 0.182 0.007 0.003 0.027
MACE 3D ID 0.000 0.084 0.156 0.024 0.017 0.000 0.000 0.000 0.049 0.001
OOD 0.145 0.035 0.007 0.012 0.004 0.000 0.083 0.004 0.035 0.002
MLP No Bias ID 0.002 0.019 0.000 0.101 0.035 0.098 0.003 0.001 0.011 0.004
OOD 0.009 0.064 0.009 0.007 0.002 0.104 0.027 0.015 0.005 0.007
MoLFormer Transformer ID 0.003 0.001 0.005 0.002 0.001 0.000 0.001 0.001 0.005 0.001
OOD 0.036 0.112 0.017 0.013 0.001 0.049 0.056 0.016 0.002 0.100
ModernBERT Transformer ID 0.012 0.003 0.029 0.006 0.012 0.000 0.009 0.006 0.030 0.002
OOD 0.061 0.005 0.058 0.076 0.121 0.002 0.000 0.190 0.015 0.011
RT Transformer ID 0.012 0.002 0.491 0.002 0.013 0.016 0.005 0.018 0.009 0.002
OOD 0.157 0.090 0.003 0.015 0.011 0.036 0.049 0.047 0.015 0.070
Random Forest No Bias ID 0.001 0.002 0.001 0.001 0.001 0.001 0.000 0.001 0.003 0.001
OOD 0.006 0.003 0.001 0.001 0.003 0.003 0.006 0.009 0.001 0.002
TGNN Explicit-Bonds ID 0.001 0.003 0.020 0.014 0.043 0.001 0.453 0.001 0.029 0.001
OOD 0.007 0.007 0.047 0.041 0.049 0.002 0.006 0.009 0.007 0.005
Table 6: Standard deviation of Batched $R^{2}$ scores of all models on OOD and ID tasks.
Model Representation Split HoF Density HOMO LUMO GAP ZPVE $\mathbf{\left<R^{2}\right>}$ $\boldsymbol{\alpha}$ $\boldsymbol{\mu}$ $\mathbf{C_{v}}$
Random SMILES ID 24.43 0.0248 0.0061 0.0071 0.0085 0.00113 52.3 0.940 0.674 0.375
Forest OOD 139.89 0.1815 0.0304 0.0372 0.0371 0.02303 363.0 8.470 2.899 3.362
MLP SMILES ID 13.66 0.0532 0.0094 0.0091 0.0130 0.00560 49.9 0.817 0.696 0.384
OOD 38.43 0.0941 0.0247 0.0201 0.0468 0.01370 470.9 6.859 2.389 0.593
RT SMILES ID 22.23 0.0163 0.0090 0.0102 0.0133 0.00289 68.2 2.264 1.104 0.654
OOD 2428764 7558.9 539.74 584.88 0.0339 27.6906 25458 69968 2.719 11435
ChemBERTa SMILES ID 22.72 0.0154 0.0070 0.0093 0.0104 0.00390 51.7 1.254 0.713 0.489
OOD 100.86 0.1195 0.0244 0.0256 0.0309 0.02253 302.6 6.328 2.723 2.765
MoLFormer SMILES ID 10.94 0.0273 0.0050 0.0052 0.0064 0.00106 40.1 1.047 0.602 0.785
OOD 93.6 0.0770 0.0236 0.0256 0.0275 0.01990 314.1 7.356 2.232 3.667
Chemprop Graph ID 15.43 0.0092 0.0041 0.0046 0.0058 0.00188 35.8 0.866 0.545 0.340
OOD 99.72 0.0347 0.0189 0.0179 0.0269 0.01277 233.7 4.850 2.304 2.118
EGNN 3D ID 10.07 0.0077 0.0048 0.0052 0.0069 0.00103 19.6 0.566 0.481 0.267
OOD 19.19 0.0279 0.0212 0.0236 0.0312 0.00583 181.3 5.659 2.446 2.079
IGNN 3D ID 14.68 0.0084 0.0050 0.0053 0.0070 0.00173 77.5 0.903 0.519 0.405
OOD 23.35 0.0281 0.0818 0.0194 0.0297 0.00677 128.6 5.611 2.501 2.212
TGNN Graph ID 14.46 0.0258 0.0057 0.0072 0.0178 0.00167 211.6 0.751 0.673 0.377
OOD 29.20 0.0331 0.0184 0.0190 0.0424 0.00260 625.5 2.787 2.524 0.627
MACE 3D ID 5.56 0.0617 0.0150 0.0135 0.0181 0.00180 9.8 0.322 0.430 0.134
OOD 38.86 0.0670 0.0339 0.0247 0.0409 0.00210 68.3 1.543 2.228 0.229
GotenNet 3D ID 5.44 0.0070 0.0052 0.0039 0.0088 0.00043 11.8 0.553 0.319 0.197
OOD 15.54 0.0360 0.0126 0.0118 0.0229 0.00053 16.7 1.825 2.173 0.302
Graphormer 3D ID 9.64 0.0068 0.0040 0.0042 0.0055 0.00024 33.4 0.431 0.626 0.180
OOD 31.62 0.0770 0.0236 0.0256 0.0275 0.01990 314.1 7.356 2.232 3.667
ET 3D ID 29.97 0.0081 0.0027 0.0031 0.0043 0.00057 28.2 0.490 0.368 0.160
OOD 52.50 0.0479 0.0220 0.0236 0.0271 0.01710 298.0 6.568 2.257 3.405
Geoformer 3D ID 17.77 0.0071 0.0027 0.0028 0.0046 0.00030 10.9 0.326 0.847 0.124
OOD 43.32 0.0366 0.0157 0.0186 0.0240 0.00557 63.8 4.201 2.544 1.354
ModernBERT SMILES ID 14.68 0.0117 0.0064 0.0076 0.0095 0.00073 40.1 0.870 0.698 0.407
OOD 44.49 0.0287 0.0216 0.0232 0.0324 0.00170 228.7 2.489 2.657 0.611
Table 7: RMSE scores of all models on OOD and ID tasks. Best performing ID and OOD models are highlighted in Black and Blue respectively.The worst performing ID and OOD models are highlighted in Orange and Red respectively. The graph-based and hybrid models provide the best scores across nearly all tasks for OOD and ID splits. Numerical encoding issues greatly hamper RTs performance and result in large errors. We additionally provide results on using a Llama large language model for OOD property prediction in Appendix E. All results are averaged across 3 training runs.
C.7 Broader Impacts
BOOM provides a set of benchmarks that are designed to accelerate the development of generalizable chemical foundation models. In turn, we aim for these chemical foundation models to be used to tackle important societal issues such as developing revolutionary pharmaceuticals or energy storage materials. Nevertheless, we note that it is important to develop appropriate safeguards to ensure that such chemistry machine learning models are not used for the development of dangerous chemicals. To this end, we advocate for the continued development of chemical safety benchmarks to assess the potential for chemistry machine learning models to design harmful materials.
C.8 Discussion of Data Augmentation
The data augmentation experiment in the main text serves to demonstrate that even a few thousand OOD samples can effectively convert an OOD region into an in-distribution (ID) one. This has important implications for real-world applicability. Our experiment shows that property prediction performance can be improved with minimal OOD data-highlighting that even a relatively small number of molecules in the OOD region can significantly enhance model generalization. Our hope with this experiment is that it inspires future work exploring how targeted generation can improve OOD generalization, rather than definitively prescribing a solution for solving OOD generalization.
Nevertheless, the feasibility of identifying useful OOD points in the first place is a significant, unsolved challenge and likely requires approaches based on Bayesian Optimization or Active Learning. Nevertheless, recent work that demonstrates that generative models combined with active learning may be able to extrapolate in property space. Prior work has shown that generative models without active learning were not able to extrapolate beyond the training data.antoniuk_active_2025 However, once active learning on DFT simulations was incorporated into the generative loop, the model showed strong potential to generate molecules with properties that extend far beyond the training data, demonstrating an ability to extrapolate in property space. In another more extreme example, the authors used their STGG+ autoregressive generative model with active learning to discover molecules with an oscillator strength of 27.7, compared to a maximum of 9.3 in their training data, and a value of 13.01 without active learning.jolicoeur-martineau_generating_2025 These two examples serve to empirically demonstrate that iteratively generating molecules, labeling them with ground-truth simulations, and then retraining the property prediction models may lead generative models to recognize extreme-property domains in chemical space.
Conceptually, we hypothesize that this approach may be possible because this iterative active learning will continually trend towards molecules with improved properties as long as the property prediction models are able to determine the relative ordering of molecules with respect to the property of interest, rather than needing quantitatively correct predictions. Then, it is the ground truth simulations (DFT) that will provide the true property labels of the molecules. Our scatter plots shown in Figure 13 as an example, show that the property prediction models do seem to have this capability, as the most extreme-property molecules are consistently identified, and thus, would be preferentially generated.
Encouragingly, the proposed overall approach of improving OOD performance through iterative active learning has already seen some reported success in the literature. In recent work, the authors note that after three iterations of active learning, the prediction RMSE reduces by $83\%$ when evaluated on hold-out test molecules from across the entire active learning run.antoniuk_active_2025 Similarly in another work, the authors found that multiple iterations of active learning to generate novel inorganic structures reduced the error on structure-based OOD energy predictions from >200meV/atom down to  25meV/atom.merchant_scaling_2023 Although there is still much to explore, we feel that there is some growing evidence to believe that OOD generalization can be improved in this manner.
C.9 Lipophilicity Dataset
To provide a further assessment of OOD performance beyond the computational datasets discussed in the main text, we also evaluate a subset of the models on the Lipophilicity Dataset from MoleculeNet.moleculenet This dataset consists of 4200 experimental measurements of the octanol/water distribution coefficient, which is of relevance for drug compounds. The inclusion of the Lipophilicity dataset serves as an exemplary dataset for performing OOD evaluations on experimentally measured properties, rather than only computed physicochemical properties.
C.10 Statistical Analysis
Figure 11: Gaussian process fitting of OOD $R^{2}$ values given ID $R^{2}$ scores. Interestingly, lower ID $R^{2}$ scores are correlated with low OOD $R^{2}$ scores, as expected. Higher ID scores are not necessarily predictive of OOD values, signalling a need to test models with ID-OOD splits as well as random splits.
Property N P-value (Kruskal-Wallis) P-value (Mann-Whitney U)
Geometric > Transformer Geometric > Graph Graph > Transformer
HoF 39 0.00000 0.00000 0.00398 0.00485
Density 39 0.00424 0.00175 0.07845 0.00673
HOMO 39 0.05119 0.01569 0.48837 0.02197
LUMO 39 0.00046 0.00010 0.20389 0.00485
GAP 39 0.10577 0.98113 0.78453 0.85461
ZPVE 39 0.05163 0.01067 0.51164 0.05052
$R^{2}$ 39 0.00006 0.00018 0.00002 0.30314
$\alpha$ 39 0.01024 0.00165 0.23808 0.05123
$\mu$ 39 0.49232 0.83475 0.76714 0.87963
$C_{v}$ 39 0.02554 0.00641 0.45356 0.02074
Table 8: We perform statistical analysis on the OOD performance of model groups while controlling for the property. We first use the Kruskal-Wallis test to detect whether there is a statistically significant difference between Geometric, Transformer, and Graph models, given a property. Then we perform the Mann-Whitney U hypothesis tests to identify orderings within the groups. Interestingly, we only fail to reject the null hypothesis for $\mu$ and GAP, as many of the models performed poorly on the two tasks.
Model MLP Random Forest Regression Transformer MoLFormer Chemprop
ID $0.866\pm 0.09$ $0.548\pm 0.001$ $1.139\pm 0.02$ $0.473\pm 0.006$ $0.463\pm 0.01$
OOD $2.041\pm 0.2$ $576\pm 0.006 1.1$$164\pm 0.003$ $0.956\pm 0.004$ $1.051\pm 0.02$
Table 9: RMSE values of various models on the Lipophilicity Dataset from MoleculeNet. We report the RMSE values, averaged across 3 training runs, along with their standard deviations.
Appendix D Parity Plots
As there are more than 150 plots, we provide the parity plots for all of our experiments in a compressed layout in the following section, intended for observing the prediction trends for the model. We also upload the higher resolution images, as well as the actual predictions, including the training/fine-tuning code for all models, to our repository.
@@ Appendix D Parity Plots
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 12: Parity Plots for Random Forest on 10K and QM9 OOD tasks.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 13: Parity Plots for Chemprop on 10K and QM9 OOD tasks.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 14: Parity Plots for Regression Transformer (with Pretraining) on 10K and QM9 OOD tasks.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 15: Parity Plots for MoLFormer (with Pretraining) on 10K and QM9 OOD tasks.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 16: Parity Plots for ChemBERTa (with Pretraining) on 10K and QM9 OOD tasks.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 17: Parity Plots for ModernBERT on 10K and QM9 OOD tasks.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 18: Parity Plots for EGNN on 10K and QM9 OOD tasks.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 19: Parity Plots for IGNN on 10K and QM9 OOD tasks.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 20: Parity Plots for TGNN on 10K and QM9 OOD tasks.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 21: Parity Plots for Graphormer(3D) on 10K and QM9 OOD tasks.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 22: Parity Plots for TorchMD-ET on 10K and QM9 OOD tasks.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 23: Parity Plots for MACE on 10K and QM9 OOD tasks.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 24: Parity Plots for GeoFormer on 10K and QM9 OOD tasks.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 25: Parity Plots for GotenNet on 10K and QM9 OOD tasks.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 26: Parity Plots for LLAMA on 10K and QM9 OOD tasks.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 27: Parity Plots for MLP on 10K and QM9 OOD tasks.
Appendix E Property Prediction with LLMs
Large language models (LLMs) have seen increasing usage for a wide range of molecular design tasks including property prediction,[jacobs_regression_2024, jablonka_is_2023] molecular synthesis prediction, and property-guided molecule design.[jablonka_is_2023, bhattacharya_large_2024] In this section, we benchmark the performance of LLMs on BOOM’s OOD property prediction tasks. We note that in the context of LLMs, it is more challenging to robustly define OOD molecules without knowing the exact training corpus of the LLM. For example, we anticipate that it is possible that the density of some of the molecules in our 10k Density OOD test set appear as natural language in the training corpus of the LLM. Although we believe that including benchmarking of the OOD performance of LLMs is important due to their widespread usage, we caution against direct comparison of the models in the main text due to these possible data leakage concerns.
E.1 Experiment Details
We use the LLAMA-3.1-8B model provided by Meta.[grattafiori_llama_2024] We use the following prompt to generate the properties:
[…]
Where, <smiles> is the SMILES representation of the molecules. <property> is one of the ten properties mentioned above. <property_description> is the description of the property, and <units> is the units of the properties.
Property Description Text Unit Text
HoF solid heat of formation (using a group additivity approach) g/$c$
Density crystalline density kCal/mol
$\alpha$ isotropic polarizability a_0^3
$C_{v}$ heat capacity at 298.15 K cal/molK
HOMO energy of the highest occupied molecular orbital Hartree energy
LUMO energy of the lowest unoccupied molecular orbital Hartree energy
Gap energy difference between the highest occupied and lowest unoccupied molecular orbital Hartree energy
$\mu$ dipole moment Debye
$\left<R^{2}\right$> electronic spatial extent a_0^3
ZPVE zero point vibrational energy Hartree energy
Table 10: Property descriptions and units used in the LLAMA prompt
We prompt the model to output only floating-point values and use a parser to extract the first decimal numerical values from the generated output.
E.2 Llama Results
Model Split HoF Density HOMO LUMO GAP ZPVE $\mathbf{\left<R^{2}\right>}$ $\boldsymbol{\alpha}$ $\boldsymbol{\mu}$ $\mathbf{C_{v}}$
LLaMA 3.1 8B ID 93.3778 0.1446 93.3778 .3097 1.0234 3.8958 1176 75.44 1.4703 57.796
(no finetuning) OOD 262.9860 0.1744 262.9860 1.1436 0.9545 4.9597 1688 73.34 5.6357 32.026
Table 11: RMSE of LLaMA models on all OOD and ID tasks.
Model Split HoF Density HOMO LUMO GAP ZPVE $\mathbf{\left<R^{2}\right>}$ $\boldsymbol{\alpha}$ $\boldsymbol{\mu}$ $\mathbf{C_{v}}$
LLaMA 3.1 8B ID 0.008 0.2050 0.0009 0.0025 0.0127 0.0008 0.0000 0.0034 0.0845 0.0000
(no finetuning) OOD 0.0254 0.1048 0.0007 0.0000 0.0018 0.0022 0.0005 0.0016 0.0631 0.0007
Table 12: Batched $R^{2}$ scores of LLaMA models on all OOD and ID tasks.